Calculate the mass of water produced when 1.41 g

of butane reacts with excess oxygen.


1
Expert's answer
2021-10-08T02:09:25-0400

"2C_4H_{10}+13O_2 \to 8CO_2+10H_2O"

Molar mass of butane "=2\times 58=116"

number of moles of butane "=1.41\/116=0.0122"

molar mass of water"=10\times 18=180"

mole ratio "2C_4H_{10}:10H_2O=2:10=1:5"

moles of water "=5\times 0.0122=0.061"

mass of water produced "=0.061\times 180=10.98g"